OpenStudy (anonymous):

There must be a variable term in both the numerator and denominator. (5x) (2x) (x + 4) (x – 5) (2x + 1) (3x + 5)

OpenStudy (anonymous):

@iGreen @ganeshie8

OpenStudy (anonymous):

Turn one. Flip your coin and perform the appropriate operation. Explain to the game master how to add your rational expression to the one on the correct space. Use complete sentences.

OpenStudy (anonymous):

I got tails

OpenStudy (igreen):

Right..so that means you do: \(\dfrac{x-1}{3x}\)

OpenStudy (anonymous):

That's it?

OpenStudy (igreen):

Actually it wants you do add both of them..so I guess it's: \((x - 1) + (3x)\) To simplify, add the like terms..

OpenStudy (anonymous):

4x-1?

OpenStudy (igreen):

Yep.

OpenStudy (anonymous):

That's the answer? Are you serious?!

OpenStudy (igreen):

Yes.

OpenStudy (igreen):

That's the answer for Turn 1, if you get Tails..

OpenStudy (anonymous):

It says I have to explain to the game master how to add your rational expression to the one on the correct space.

OpenStudy (igreen):

Huh..?

OpenStudy (igreen):

Ohhhh..you have to make a rational expression..then you add it with \(\dfrac{x-1}{3x}\).

OpenStudy (anonymous):

But what do I do with the given examples (5x) (2x) (x + 4) (x – 5) (2x + 1) (3x + 5)

OpenStudy (igreen):

It says there has to be a numerator in the top and bottom..so make one out of the one's they gave you.

OpenStudy (igreen):

Just have '5x' or '2x' be the denominator, and the others the numerator.

OpenStudy (anonymous):

so it would be \[\frac{ 4x-1 }{ 2x }\]

OpenStudy (igreen):

No..you have to pick a numerator out of: (x + 4) (x - 5) (2x + 1) (3x + 5)

OpenStudy (igreen):

Which one do you want to pick?

OpenStudy (anonymous):

the third, but why can't I have 2x or 5x be the denominator?

OpenStudy (igreen):

You can..but you just can't have \(4x - 1\) as the numerator..

OpenStudy (igreen):

So therefore our expression is: \(\dfrac{2x + 1}{2x}\) Now since you got Tails, we can add this to \(\dfrac{x-1}{3x}\) So can you add this? \(\dfrac{2x + 1}{2x} + \dfrac{x - 1}{3x}\)

OpenStudy (anonymous):

\[\frac{ 8x+1 }{ 6x }\]?

OpenStudy (igreen):

You got it.

OpenStudy (igreen):

Now flip the coin again.

OpenStudy (anonymous):

Tails again

OpenStudy (anonymous):

Turn two. Flip your coin and perform the appropriate operation. Discuss and identify any possible restrictions that exist with (or in) the resulting rational expression.

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

Okay, now we have to subtract. \(\dfrac{2x + 1}{2x} - \dfrac{x}{x-4}\)

OpenStudy (igreen):

Can you subtract that?

OpenStudy (anonymous):

\[-\frac{ 7x+4 }{ 2x(x-4) }?\]

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

Yep.

OpenStudy (igreen):

Now flip the coin again.

OpenStudy (anonymous):

I need to find any restrictions that exist with, or in, the resulting expression before I can move on.

OpenStudy (igreen):

What do you mean by restrictions?

OpenStudy (anonymous):

Anything that could prevent the following outcome from presenting itself.

OpenStudy (igreen):

Hmm..I don't see any of those :l

OpenStudy (anonymous):

So I'll just put 'No identified restrictions'

OpenStudy (anonymous):

Heads: Turn three. Flip your coin and perform the appropriate operation. Explain to the game master how to multiply your rational expression to the one on the correct space. Use complete sentences.

OpenStudy (igreen):

Okay, so now we multiply: \(\dfrac{2x + 1}{2x} \times \dfrac{x-4}{x+1}\) What do you get?

OpenStudy (anonymous):

\[\frac{ (2x+1)(x-4) }{ 2x(x+1) }\]

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

Yes..and flip the coin again..

OpenStudy (anonymous):

Tails: Turn four. Flip your coin and perform the appropriate operation. Discuss why the degree of the resulting denominator did not change from your expression’s degree.

OpenStudy (igreen):

Okay, so we divide: \(\dfrac{2x + 1}{2x} \div \dfrac{6}{x-1}\)

OpenStudy (anonymous):

\[\frac{ (2x+1)(x-1) }{ 12x }\]

OpenStudy (igreen):

Yep..now we move onto the final step..

OpenStudy (igreen):

I guess we add: \(\dfrac{2x+1}{2x} + \dfrac{x}{x+1} + \dfrac{x}{x-1}\)

OpenStudy (anonymous):

Before we can move on, I need to discover why the degree of the resulting denominator did not change from my expression's degree.

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

Because when you divided it you got the same degree..

OpenStudy (igreen):

(Which is a degree of 1)

OpenStudy (anonymous):

You put dots at the end as if it is a stupid question lol

OpenStudy (anonymous):

\[\frac{ 6x^3+x^2-2x-1 }{ 2x(x+1)(x-1) }\]

OpenStudy (anonymous):

for the last one ^^^

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

Yep, you got it.!

OpenStudy (anonymous):

Thanks so much!

OpenStudy (igreen):

No problem, you did pretty much all the work :P

OpenStudy (anonymous):

You were very patient with my incompetence though.